Determine the molality of an aqueous solution prepared by dissolving 0.224 moles of LiBr in 27.4 moles of H2O.
Answer #1

Molaloty = number of moles of solute/ kg of solvent

molality of aqueous solution of LiBr= moles of LiBr/ kg of water

moles of LiBr= 0.224 moles, Moles of water= 27.4, mass of water= Moles* Molecular weight= 27.4*18=493.2 gms

=0.4932 Kgs

Molality= 0.224/0.4932 =0.4642